Two cars started moving with initial velocities u and 2u. For the same deceleration, their respective stopping distances are in the ratio

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
DEAR STUDENT

Explanation:

using V^2 = U^2 - 2*a*s

s = v^2 / (2*a)

their respective stopping distances are in the ratio

1:4

THANKS